Natalie is thinking of buying a van that will be used only for business. The cost of the van is estimated at $36,500. Natalie would spend an additional $2,500 to have the van painted. In addition, she wants the back seat of the van removed so that she will have lots of room to transport her mixer inventory as well as her baking supplies. The cost of taking out the back seat and installing shelving units is estimated at $1,500. She expects the van to last about 5 years, and she expects to drive it for 200,000 miles. The annual cost of vehicle insurance will be $2,400. Natalie estimates that at the end of the 5-year useful life the van will sell for $7,500. Assume that she will buy the van on August 15, 2017, and it will be ready for use on September 1, 2017. Natalie is concerned about the impact of the van’s cost on her income statement and balance sheet. She has come to you for advice on calculating the van’s depreciation. Prepare three depreciation tables for 2017, 2018 and 2019: one for straight-line depreciation , one for double-declining balance depreciation, and one for units-of-activity depreciation. For units-of-activity, Natalie estimates she will drive the van as follows: 15,000 miles in 2017; 45,000 miles in 2018; 50,000 miles in 2019; 50,000 miles in 2020; and 40,000 miles in 2021. Recall that Cookie Creations has a December 31 year-end.

Double-declining-balance depreciation: What is just the NBV (Beg. of Year) for 2017, 2018, and 2019?

Double-declining-balance depreciation

NBV(BEG OF YEAR) depreciation exp    accumulated dep netbook value   
2017 ??????
2018 ??????
2019 ?????

Units-of activity depreciation:What are the Units of Activity for 2017, 2018, and 2019?

Units of activiy depreciation

Units of activity Depreciation Exp Accumulated Deprec Net book Value
2017 ????
2018 ????
2019 ?????
Answer #1

Costs to be considered for depreciation are as follows:

Particulars Amount
Cost of the van $36,500.00
Cost of painting the Van $2,500.00
Cost of installing shelving units $1,500.00
Total cost of the van $40,500.00

Answer for question no.1:

Depreciation under straight line basis = ($40,500 -$7,500)/5

=$33,000/5

=$6,600.

As the van is in operation only for seven months in the first year, depreciation is to be calculated for 7 months

=$6,600*7/12

=$3,850.

Answer for question no.2:

Given salvage value after 5 years

Depreciation rate=1/5=20%

As it is double declining rate of depreciation=20%*2=40%.

Under mentioned is the schedule fore depreciation.

Year NBV (Beginning of the year)(A) Depreciation expense(B)=(A)*40% Accumulated depreciation(C )=Current year depreciation +previous years depreciation Net book value=(A)-(B)
2017 $40,500.00 $4,725.00 $4,725.00 $35,775.00
2018 $35,775.00 $14,310.00 $19,035.00 $16,740.00
2019 $16,740.00 $6,696.00 $25,731.00 $10,044.00

As in the first year of operation, the van is put to use only for seven month so, depreciation is to be calculated for 7 month only ie., $40,500*40%*7/12

Answer for question no.3:

Cost of machine - salvage value=$40,500 -$7,500=$33,000.

Total units of activity over the life of the van=200,000 miles.

Depreciation per mile=$33,000./200000

=$0.17

Depreciation schedule as required is below:

Year Units of activity(A) Depreciation expense(B)=(A)*0.17 Accumulated depreciation(C )=Current year depreciation +previous years depreciation Net book value=40500-(C )
2017 15000 $2,550.00 $2,550.00 $37,950.00
2018 45000 $7,650.00 $10,200.00 $30,300.00
2019 50000 $8,500.00 $18,700.00 $21,800.00